Harry went to the electronic shop to buy a 17.5 ohm resistor for his project. However, the store only has 5 ohm resistor. What is the minimum number of 5 ohm resistors must he buy? How will he connect these resistors to come up with 17.5 ohm resistance, he is expected to use all the resistors that he will buy.


1
Expert's answer
2021-05-03T10:34:03-0400

"17.5+5\cdot3=2.5 \ (\Omega)"


"n_1=3"


"2.5=1/(n_2/5)\to n_2=2"


"n=n_1+n_2=3+2=5" . Answer


Three resistors must be connected in series and two in parallel.


"5+5+5+(\frac{1}{1/5+1/5})=17.5\ (\Omega)"